2. A laboratory group measures an excess of an unknown liquid. They add this to an empty flask for which they calculate the volume. They heat the sample in a water bath until all liquid has been vaporized for several minutes. The sample is cooled and measured. Use their laboratory data (isted below) to answer the following questions Flask Volume 100.0 100.0 75.6 mL 275.6 m L 145.81 g Mass of flask 149.25 g Mass of Flask volatile liquid 3.44 g Mass of volatile liquid 84.1 C Temperature of water bath (C) 0.914 atm Pressure in lab (atm) 0.0821 (L atm) /(mol K) Gas Constant a. Calculate the mol of gas in the flask using the ideal gas law: (o quy ain) 0.2756 L 0,01 mol mil W b. Calculate the molar mass of the unknown: Pase answe 3. If a student added more than 4 mL of volatile liquid, would it affect the laboratory today? How? What if less liquid was added?
Answer #1

a)

volume of flask = 275.6 mL = 0.2756 L

pressure = 0.914 atm

temperature = 84.1 + 273.15 = 357.25 K

P V = n R T

0.914 x 0.2756 = n x 0.0821 x 357.25

n = 8.588 x 10^-3

moles of gas in the flask = 8.588 x 10^-3

b)

mass of volatile liquid = 149.25 - 145.81 = 3.44 g

moles = mass / molar mass

8.588 x 10^-3 = 3.44 / molar mass

molar mass of unknown = 400.6 g/mol